# Dirichlet process mixtures of block $g$ priors for model selection and prediction in linear models

### v1

@@ 3 A new class of priors: Dirichlet process mixtures of block $g$ priors
To conclude this section, we note that the specification in (3.1) was proposed by Finegold & Drton (2014) in the context of Gaussian graphical models. However, Finegold & Drton (2014) used (3.1) to model the observed data rather than as a prior distribution for unknown parameters. Furthermore, their key variable of interest was the matrix $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}$, which is fixed in our case.
3.1 A unifying framework for continuous shrinkage and variable selection priors
Hierarchical priors based on (3.1) provide a unifying framework for thinking about continuous shrinkage priors and model selection priors. To see this, consider a slightly less general spike-and-slab version of our prior where
$\boldsymbol{\beta}\mid\sigma^{2},\tilde{g}_{1},\ldots,\tilde{g}_{p},\boldsymbol{\gamma}\sim\mathsf{N}\left(\boldsymbol{0},\sigma^{2}\tilde{\boldsymbol{G}}^{1/2}\left\{\boldsymbol{\Gamma}\boldsymbol{\Sigma}^{-1}\boldsymbol{\Gamma}\right\}^{-}\tilde{\boldsymbol{G}}^{1/2}\right),$
$\boldsymbol{\Gamma}=\operatorname{diag}\{\gamma_{1},\ldots,\gamma_{p}\}$, $\tilde{\boldsymbol{G}}=\operatorname{diag}\{\tilde{g}_{\xi_{1}},\ldots,\tilde{g}_{\xi_{p}}\}$, $\tilde{g}_{1},\tilde{g}_{2},\ldots$ are i.i.d. from (3.3), $\xi_{1},\ldots,\xi_{p}$ are i.i.d. from a categorical distribution, and $A^{-}$ represents the Moore-Penrose inverse of $A$. Different choices of $\boldsymbol{\Sigma}$ and of priors on $\boldsymbol{\gamma}$ and $\rho$ lead to various well-known procedures. For example, as we noted before, the DP mixture of block $g$ priors includes the standard $g$ prior as a special (limit) case. Furthermore, when $\boldsymbol{X}$ is orthogonal and the grouping variable $\rho$ is treated as known, it also includes the block $g$ prior of Som (2014) as a special case. However, our procedure is not equivalent to that of Som (2014) in the more general case where the $\boldsymbol{X}$ is not orthogonal.
@@ 3.1 A unifying framework for continuous shrinkage and variable selection priors
On the other hand, if we fix the model to $\boldsymbol{\gamma}=(1,1,\ldots,1)$, either $\boldsymbol{X}$ is orthonormal or we take $\boldsymbol{\Sigma}=\boldsymbol{I}$, and $\tau^{2}$ is given a hyperprior, the DP mixture of block $g$ priors corresponds to the Horseshoe Pit mixture prior of Denti et al. (2023). As a consequence, it also includes a number of traditional continuous shrinkage priors. For example, when the mixing density, $f(g)$ is exponential, we recover the Bayesian LASSO (Park & Casella, 2008; Hans, 2009). Similarly, when the mixing density is half-Cauchy, then we recover the Horseshoe prior (Carvalho et al., 2010). A more extensive but non-exhaustive list of examples includes the Student $t$-prior (Tipping, 2001), Normal-Exponential-Gamma prior (Griffin & Brown, 2005), Normal-Gamma prior (Brown & Griffin, 2010), the Dirichlet-Laplace prior (Bhattacharya et al., 2015), global-local shrinkage priors (Polson & Scott, 2012), Horseshoe$+$ prior, (Bhadra et al., 2017), Beta-prime prior (Bai & Ghosh, 2018), and tail-adaptive shrinkage prior (Lee et al., 2020). Also under the full model, and under the block orthogonality and known group structure assumptions, the DP mixture of block $g$ priors includes the Group Inverse Gamma Gamma shrinkage prior of Boss et al. (2023) as a special case.
The discussion above highlights that DP mixtures of block $g$ priors provide a unified framework for thinking about model selection and continuous shrinkage priors, providing a mechanism to address some of the shortcomings of both types of approaches. Traditionally, frameworks based on continuous shrinkage priors have recognized that differential shrinkage might be needed to attain optimal performance, but until recently (e.g., Boss et al., 2023) they have tended to downplay the need to account for co-linearity among covariates. On the other hand, the literature on priors for model selection has, from the very beginning, acknowledged the need to account for colinearity, but has been slower to recognize the need for differential shrinkage, perhaps because of the computational challenges involved. Dirichlet process mixtures of block $g$ priors incorporate the best of both worlds, allowing for differential shrinkage among variables in a given model as well as for full model uncertainty quantification.
4 Properties of the DP mixtures of block $g$ prior
@@ 4.1 Tail behavior
Because, marginally, samples from a distribution generated by a Dirichlet process follow the baseline measure (e.g., see Blackwell & MacQueen, 1973 or Antoniak, 1974), the marginal distribution for the $l$-th entry of $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ under a DP mixture of block $g$ priors is given by
$f(\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\mid\tau^{2},a,b,\gamma)=\int\mathsf{N}(\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\mid 0,g\kappa_{\gamma,l,l}\sigma^{2})f(g\mid\tau^{2},a,b)\mathop{}\!\mathrm{d}g$
where $\kappa_{\gamma,l,l}$ is the $l$-th diagonal entry of $\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$.The tail behavior of this type of marginal distributions was studied in Boss et al. (2023) (see their Theorem 2.1). In particular, the index of regular variation of the marginal prior is $\omega=-2b-3$, i.e.,
@@
This implies that the our prior has heavy (polynomial) tails and any point estimator derived from our procedures are robust, in the sense of having bounded influence in the case of likelihood-prior conflict (Andrade & O’Hagan, 2011). However, note that, while the marginal tail behavior of the DP mixture of block $g$ priors is identical to that of the regular $g$ prior that uses the same Beta-prime distribution for the common shrinkage factor, the contour plots associated with the DP mixture of block $g$ priors are not elliptical. Figure 2 illustrates this in the bivariate case.
(a) Standard $g$ prior
(b) Orthogonal block $g$ prior
(c) Block $g$ prior
(d) DP mixture of block $g$ priors
Figure 2: Scatterplots of random samples from the Dirchlet mixture of block $g$ priors and some related distributions in the bivariate case under a hyper-$g/n$ distributon for the shrinkage parameter(s). Panel (a) corresponds to the (elliptical) contours of the standard $g$ prior of Liang et al. (2008). Panel (b) shows the density of the prior proposed by Som et al. (2016), which assumes that blocks are orthogonal a priori. Panel (c) corresponds to a global-local $g$ prior in which each covariate is assigned is own independent shrinkage factor and the prior covariance matrix is proportional to $\left(\boldsymbol{X}^{T}\boldsymbol{X}\right)^{-1}$. Panel (d) is our DP mixture of block $g$ priors, which in this case corresponds to a mixture of the distributions in panels (a) and (c).
@@ 4.2 Information consistency of Bayes factors
For fixed $n$, $p$ and $\boldsymbol{X}$ that is full rank, consider a sequence of observations $\boldsymbol{y}(1),\boldsymbol{y}(2),\ldots$ such that $\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}(N)\|\to\infty$ and $N\to\infty$, where $\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}(N)=\left[\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right]^{-1}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{y}(N)$ is the maximum likelihood estimator or $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ based on $\boldsymbol{y}(N)$. The Bayes factor $BF_{\boldsymbol{\gamma},\boldsymbol{0}}$ is information consistent if $BF_{\boldsymbol{\gamma},\boldsymbol{0}}\left(\boldsymbol{y}(N)\right)\to\infty$ as $N\to\infty$. Intuitively, this means that, as the information in the likelihood favors model $\boldsymbol{\gamma}$ over the null model, the Bayes factor grows without bounds.
@@
The proof is included in Section B of the supplementary materials. It is worthwhile noting that the integrand in (4.1) has tails that behave as polynomials of $g_{1}^{1/2},\ldots,g_{p_{\boldsymbol{\gamma}}}^{1/2}$, with the order of the polynomial in each dimension being a function of the multiplicity of the respective $g_{j}$. A slightly simpler condition that applies to DP mixtures of block $g$ priors is the following:
@@
A sufficient condition for the DP mixture of block $g$ priors to lead to Bayes factors that are information consistent is for the density of the centering measure, $f(g\mid\tau^{2},a,b)$, to satisfy
$\int(1+g_{j})^{(n-1-p_{\boldsymbol{\gamma}})/2}f(g_{j}\mid\tau^{2},a,b)\mathop{}\!\mathrm{d}g_{j}=\infty$
for all $p_{\boldsymbol{\gamma}}\leq p$.
See Section B of the supplementary materials. This is the same condition in Theorem 2 of Liang et al. (2008). Hence, this result just indicates that any mixing distribution for $g$ that leads to information-consistent Bayes factors under a standard mixture of $g$ priors also leads to information consistent Bayes factors under the DP mixture of block $g$ priors.
@@ 4.3 Information consistency of block structures
One of the key motivations to consider DP mixture of block $g$ priors is the desire to avoid having to decide a priori on an appropriate partition for the covariates. In this Section, we show that, when the design matrix $\boldsymbol{X}$ is orthogonal and the true coefficients have very different sizes, our prior assigns coefficients of different sizes to different clusters with high probability. Before proceeding with our main result, we need to introduce the concept of refinement of a partition (sometimes called a fragmentation, e.g., see Bertoin, 2006).
@@
Let $\rho=\{S_{1},\dots,S_{K}\}$ and $\rho^{\prime}=\{S_{1}^{\prime},\dots,S_{K^{\prime}}^{{}^{\prime}}\}$ denote two partition of a set =$\mathcal{I}=\{1,\ldots,c\}$ with $K$ and $K^{\prime}$ unique blocks respectively, such that $1\leq K^{\prime}\leq K\leq c$. Then, $\rho$ is said to be a refinement of $\rho^{\prime}$, denoted by $\rho\prec\rho^{\prime}$, if and only if for every $S_{k}\in\rho$ there exist a $S_{j}^{{}^{\prime}}\in\rho^{\prime}$ such that $S_{k}\subseteq S_{j}^{{}^{\prime}}$.
@@
Let $\boldsymbol{X}$ be a full rank, centered, orthogonal design matrix of size $n\times p$, and $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$ be two non-overlapping submatrices of sizes $n\times p_{1}$ and $n\times p_{2}$ with $p_{1}>0$, $p_{2}>0$ and $p_{1}+p_{2}\leq p$. Denote by $\mathcal{I}_{1}=\{j^{(1)}_{1},\ldots,j^{(1)}_{p_{1}}\}$ the set of indexes associated with the columns of $\boldsymbol{X}$ included in $\boldsymbol{X}_{1}$ and $\mathcal{I}_{2}=\{j^{(2)}_{1},\ldots,j^{(2)}_{p_{2}}\}$ the columns associated with $\boldsymbol{X}_{2}$, so that $\mathcal{I}=\mathcal{I}_{1}\cup\mathcal{I}_{2}$ and $\mathcal{I}_{1}\cap\mathcal{I}_{2}=\emptyset$.
Consider an asymptotic regime where, for fixed $n$, $p_{1}$, $p_{2}$, $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}$ and $\boldsymbol{\epsilon}$, a sequence of observations $\{\boldsymbol{y}(N):N\in\mathbb{N}\}$ is generated as $\boldsymbol{y}(N)=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}(N)+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon}$, where $\{\boldsymbol{\beta}_{1}(N):N\in\mathbb{N}\}$ is a sequence such that, $\beta_{j}^{2}(N)\sim\mathcal{O}(N)$ for all $j\in\mathcal{I}_{1}$. If $\rho_{0}=\{\mathcal{I}_{1},\mathcal{I}_{2}\}$, then
@@
    For $\rho\prec\rho_{0}$, $\lim_{N\to\infty}\frac{f\left(\boldsymbol{y}(N)\mid\rho\right)}{f\left(\boldsymbol{y}(N)\mid\rho_{0}\right)}=c_{\rho},$
where $0<c_{\rho}<\infty$.
The proof can be seen in Section C of the supplementary materials.
4.4 Dirichlet process mixtures of block-$g$ priors and the conditional Lindley paradox
@@
The proof of this theorem is available in Section D of the supplementary materials. Orthogonality plays a key role in the proof of this theorem, just like block orthogonality is key to similar proofs in Som (2014). While similar theoretical results are not available at this time in the non-orthogonal case, simulation studies suggest that they also hold in that case.
@@ 4.5 Model selection consistency
Model selection consistency refers to the ability of the procedure to choose the correct model as the sample size grows. DP mixtures of block $g$ priors are model selection consistent in the fixed $p$ regime (see Section E of the supplementary materials).
@@
    The smallest eigenvalue of $\boldsymbol{X}^{T}\boldsymbol{X}/n$ is bounded from below by a positive constant for all $n$.
Then, under the DP mixture block $g$ priors with $\tau^{2}=n$,
$\lim_{n\to\infty}\mathsf{Pr}(\boldsymbol{\gamma}=\boldsymbol{\gamma}_{T}\mid\boldsymbol{y})=1,$
as long as the prior on models satisfies $f(\boldsymbol{\gamma}_{T})>0$.

### v2

@@ 3 A new class of priors: Dirichlet process mixtures of block $g$ priors
3.1 Unifying continuous shrinkage and variable selection priors
Traditionally, frameworks based on continuous shrinkage priors have recognized that differential shrinkage might be needed to attain optimal performance, but until recently (e.g., Boss et al., 2023) they have tended to downplay the need to account for co-linearity among covariates. On the other hand, the literature on priors for model selection has, from the very beginning, acknowledged the need to account for colinearity, but has been slower to recognize the need for differential shrinkage, perhaps because of the computational challenges involved. Our approach provide a unifying framework for thinking about these two strands of the literature. To see this, consider a slightly less general spike-and-slab version of our prior where $\boldsymbol{\beta}\mid\sigma^{2},\tilde{g}_{1},\ldots,\tilde{g}_{p},\boldsymbol{\gamma}\sim\mathsf{N}\left(\boldsymbol{0},\sigma^{2}\tilde{\boldsymbol{G}}^{1/2}\left\{\boldsymbol{\Gamma}\boldsymbol{\Sigma}^{-1}\boldsymbol{\Gamma}\right\}^{-}\tilde{\boldsymbol{G}}^{1/2}\right)$, $\boldsymbol{\Gamma}=\operatorname{diag}\{\gamma_{1},\ldots,\gamma_{p}\}$, $\tilde{\boldsymbol{G}}=\operatorname{diag}\{\tilde{g}_{\xi_{1}},\ldots,\tilde{g}_{\xi_{p}}\}$, and $A^{-}$ represents the Moore-Penrose inverse of $A$. Different choices of $\boldsymbol{\Sigma}$ and of priors on $\boldsymbol{\gamma}$ and $\rho$ lead to various well-known procedures. For example, as we noted before, the DP mixture of block $g$ priors includes the standard $g$ prior and the “global-local” block $g$ prior as special (limit) cases. Furthermore, when $\boldsymbol{X}$ is orthogonal and the grouping variable $\rho$ is treated as known, it also includes the block $g$ prior of Som (2014) as a special case.
@@ 3.1 Unifying continuous shrinkage and variable selection priors
On the other hand, if we fix the model to $\boldsymbol{\gamma}=(1,1,\ldots,1)$, either $\boldsymbol{X}$ is orthonormal or we take $\boldsymbol{\Sigma}=\boldsymbol{I}$, and $\tau^{2}$ is given a hyperprior, the DP mixture of block $g$ priors corresponds to the Horseshoe Pit mixture prior of Denti et al. (2023). As a consequence, it also includes a number of traditional continuous shrinkage priors such as those in Park and Casella (2008), Brown and Griffin (2010), Carvalho et al. (2010), Leng et al. (2014), Bhattacharya et al. (2015) and Bai and Ghosh (2018). Also, under the full model and block orthogonality and a known group structure, our framework also includes the Group Inverse Gamma Gamma shrinkage prior of Boss et al. (2023).
4 Properties of DP mixtures of block $g$ prior
@@ 4.1 Tail behavior
Since, marginally, samples from a distribution generated by a Dirichlet process follow the baseline measure (e.g., see Blackwell and MacQueen, 1973 or Antoniak, 1974), the marginal distribution for the $l$-th entry of $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ under a DP mixture of block $g$ priors is given by $f(\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\mid\tau^{2},a,b,\gamma)=\int\mathsf{N}(\boldsymbol{\beta}_{\boldsymbol{\gamma},l}\mid 0,g\kappa_{\gamma,l,l}\sigma^{2})f(g\mid\tau^{2},a,b)\mathop{}\!\mathrm{d}g$, where $\kappa_{\gamma,l,l}$ is the $l$-th diagonal entry of $\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$. The tail behavior of this type of marginal distributions was studied in Boss et al. (2023) (see their Theorem 2.1). In particular, the index of regular variation of the marginal prior is $\omega=-2b-3$, i.e.,
@@
This implies that the our prior has heavy (polynomial) tails and point estimators derived from our procedures is robust, in the sense of having bounded influence in the case of likelihood-prior conflict. However, note that the contour plots associated with the DP mixture of block $g$ priors are not elliptical, unlike those of the standard $g$ prior (please see Section LABEL:ap:scatterplotsbivariagte of the supplementary materials). These non-elliptical contours enable non-uniform shrinkage by the DP block $g$ priors, which in turn is what enables them to skirt the conditional Lindley paradox.
@@ 4.2 Information consistency of Bayes factors
For fixed $n$, $p$ and $\boldsymbol{X}$ that is full rank, consider a sequence of observations $\boldsymbol{y}(1),\boldsymbol{y}(2),\ldots$ such that $\|\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}(N)\|\to\infty$ and $N\to\infty$, where $\hat{\boldsymbol{\beta}}_{\boldsymbol{\gamma}}(N)=\left[\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right]^{-1}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{y}(N)$ is the maximum likelihood estimator or $\boldsymbol{\beta}_{\boldsymbol{\gamma}}$ based on $\boldsymbol{y}(N)$. The Bayes factor $BF_{\boldsymbol{\gamma},\boldsymbol{0}}$ is information consistent if $BF_{\boldsymbol{\gamma},\boldsymbol{0}}\left(\boldsymbol{y}(N)\right)\to\infty$ as $N\to\infty$.
@@
The proof is included in Section LABEL:ap:informationconsistency of the supplementary materials. A slightly simpler condition that applies to DP mixtures of block $g$ priors is the following:
@@
A sufficient condition for the DP mixture of block $g$ priors to lead to Bayes factors that are information consistent is for the density of the centering measure, $f(g\mid\tau^{2},a,b)$, to satisfy $\int(1+g_{j})^{(n-1-p_{\boldsymbol{\gamma}})/2}f(g_{j}\mid\tau^{2},a,b)\mathop{}\!\mathrm{d}g_{j}=\infty$ for all $p_{\boldsymbol{\gamma}}\leq p$.
See Section LABEL:ap:informationconsistency of the supplementary materials. This is the same condition in Theorem 2 of Liang et al. (2008). Hence, this result just indicates that any mixing distribution for $g$ that leads to information-consistent Bayes factors under a standard mixture of $g$ priors also leads to information consistent Bayes factors under the DP mixture of block $g$ priors.
@@ 4.3 Information consistency of block structures
One of the key motivations to consider DP mixtures of block $g$ priors is the desire to avoid having to decide a priori on an appropriate partition for the covariates. In this Section, we show that, when the design matrix $\boldsymbol{X}$ is orthogonal and the true coefficients have very different sizes, our prior assigns coefficients of different sizes to separate clusters with high probability. Before proceeding with our main result, we need to introduce the concept of refinement of a partition (sometimes called a fragmentation, e.g., see Bertoin, 2006).
@@
Let $\rho=\{S_{1},\dots,S_{K}\}$ and $\rho^{\prime}=\{S_{1}^{\prime},\dots,S_{K^{\prime}}^{{}^{\prime}}\}$ denote two partition of a set $\mathcal{I}=\{1,\ldots,c\}$ with $K$ and $K^{\prime}$ unique blocks respectively, such that $1\leq K^{\prime}\leq K\leq c$. Then, $\rho$ is said to be a refinement of $\rho^{\prime}$, denoted by $\rho\prec\rho^{\prime}$, if and only if for every $S_{k}\in\rho$ there exist a $S_{j}^{{}^{\prime}}\in\rho^{\prime}$ such that $S_{k}\subseteq S_{j}^{{}^{\prime}}$.
@@
Let $\boldsymbol{X}$ be a full rank, centered, orthogonal design matrix of size $n\times p$, and $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$ be two non-overlapping submatrices of sizes $n\times p_{1}$ and $n\times p_{2}$ with $p_{1}>0$, $p_{2}>0$ and $p_{1}+p_{2}\leq p$. Denote by $\mathcal{I}_{1}=\{j^{(1)}_{1},\ldots,j^{(1)}_{p_{1}}\}$ the set of indexes associated with the columns of $\boldsymbol{X}$ included in $\boldsymbol{X}_{1}$ and $\mathcal{I}_{2}=\{j^{(2)}_{1},\ldots,j^{(2)}_{p_{2}}\}$ the columns associated with $\boldsymbol{X}_{2}$, so that $\mathcal{I}=\mathcal{I}_{1}\cup\mathcal{I}_{2}$ and $\mathcal{I}_{1}\cap\mathcal{I}_{2}=\emptyset$. Consider now an asymptotic regime where, for fixed $n$, $p_{1}$, $p_{2}$, $\boldsymbol{X}_{1}$, $\boldsymbol{X}_{2}$, $\beta_{0}$, $\boldsymbol{\beta}_{2}$ and $\boldsymbol{\epsilon}$, a sequence of observations $\{\boldsymbol{y}(N):N\in\mathbb{N}\}$ is generated as $\boldsymbol{y}(N)=\boldsymbol{1}_{n}\beta_{0}+\boldsymbol{X}_{1}\boldsymbol{\beta}_{1}(N)+\boldsymbol{X}_{2}\boldsymbol{\beta}_{2}+\boldsymbol{\epsilon}$, where $\{\boldsymbol{\beta}_{1}(N):N\in\mathbb{N}\}$ is a sequence such that, $\beta_{j}^{2}(N)\sim\mathcal{O}(N)$ for all $j\in\mathcal{I}_{1}$. If $\rho_{0}=\{\mathcal{I}_{1},\mathcal{I}_{2}\}$, then
@@
    For $\rho\prec\rho_{0}$, $\lim_{N\to\infty}\frac{f\left(\boldsymbol{y}(N)\mid\rho\right)}{f\left(\boldsymbol{y}(N)\mid\rho_{0}\right)}=c_{\rho}$, with $0<c_{\rho}<\infty$.
The proof can be seen in Section LABEL:ap:cluseringprop of the supplementary materials.
4.4 Conditional Lindley paradox
@@
The proof, which can be found in Section LABEL:ap:condLindleyparadox of the supplementary materials, only requires that the prior on the partition $\rho$ puts positive probability on at least one refinement of the true partition $\rho_{0}=\{\mathcal{I}_{1},\mathcal{I}_{2}\}$. This is trivially true for the Dirichlet process, whether its concentration parameter is fixed or treated as unknown and given a prior distribution. In fact, this is a very mild requirement that is also satisfied by the Bayes factor constructed under the global-local $g$ prior, and one that does not require that the partition model be able to identify the true partition $\rho_{0}$. On the other hand, orthogonality plays a key role in the proof of this theorem, just like block orthogonality is key to similar proofs in Som (2014). Simulation studies suggest that the result hold in the non-orthogonal case.
@@ 4.5 Model selection consistency
Model selection consistency refers to the ability of the procedure to choose the correct model as the sample size grows. DP mixtures of block $g$ priors are model selection consistent in the fixed $p$ regime (see Section LABEL:ap:modelselconsistency of the supplementary materials).
@@
    The smallest eigenvalue of $\boldsymbol{X}^{T}\boldsymbol{X}/n$ is lower bounded by a positive constant for all $n$.
Then, under the DP mixture block $g$ priors with $\tau^{2}=n$, $\lim_{n\to\infty}\mathsf{Pr}(\boldsymbol{\gamma}=\boldsymbol{\gamma}_{T}\mid\boldsymbol{y})=1$, as long as the prior on models satisfies $f(\boldsymbol{\gamma}_{T})>0$.